1. Consider the following table two ice cream firms output per hour. Pints of Ice Cream 30 Cones Sweet Peaks 20or Cold Stone 15 30 or . If Sweet Peaks allocated all of its time to producing pints of ice cream, how many pints could they produce in 1 hour? How many cones? (1 point) Who has absolute advantage in producing cones? Who has absolute advantage in producing pints of ice cream? (1 point) b. c Fill out the following table:(2 points) thint: think of the opportxnity cost o 1 cone as the membeγ。 pints ofice cream afirm forgoes producing because they are producing that 1 cone. Remember, the answers to these will oft be fractions/decimals Sweet Peaks Cold Stone Opportunity Cost of Producing Cone or 1 Cone o 1 Cone Opportunity Cost of Producingir 1 Pint o 1 Pint of Ice Cream Based on their respective opportunities costs, who has comparative advantage in producing cones? Who has comparative advantage in producing ice cream pints? (2 points) d.
Answer #1

Ans

1 30 and 20 respectively

2 sweat peaks in cones because it can produce more of them and none in ice cream since both produce equally

3 30/20= 1.5 30/15=2

20/30=0.67 15/30=0.5

4 sweat peaks in cones because of lower opportunity cost. Cold stone in ice creams